# The T2K Collaboration Results from the T2K experiment on neutrino mixing including a new far detector -like sample K. Abe

### v2

@@
††thanks: also at Université Paris-Saclay††thanks: also at Departament de Fisica de la Universitat Autonoma de Barcelona, Barcelona, Spain††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Moscow Institute of Physics and Technology (MIPT), Moscow region, Russia and National Research Nuclear University “MEPhI”, Moscow, Russia††thanks: also at IPSA-DRII, France††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at JINR, Dubna, Russia††thanks: also at the Graduate University of Science and Technology, Vietnam Academy of Science and Technology††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at BMCC/CUNY, Science Department, New York, New York, U.S.A.††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan
@@ D. Ferlewicz
University of Tokyo, Department of Physics, Tokyo, Japan
@@ A. Muñoz
University of Tokyo, Department of Physics, Tokyo, Japan
@@ Abstract
T2K has made improved measurements of three-flavor neutrino mixing with 19.7(16.3) protons on target in (anti-)neutrino-enhanced beam modes. A new sample of muon-neutrino events with tagged pions has been added at the far detector, increasing the neutrino-enhanced muon-neutrino sample size by 42.5%. In addition, new samples have been added at the near detector, and significant improvements have been made to the flux and neutrino interaction modeling. T2K data continues to prefer the normal mass ordering and upper octant of with a near-maximal value of the charge-parity violating phase with best-fit values in the normal ordering of , and eV2.
@@
Introduction— The T2K experiment [1] measures three-flavor neutrino mixing parameters by observing the disappearance of () and the appearance of () over a distance of 295 km in a narrow-band, predominantly or beam, which peaks at an energy of 0.6 GeV. Data corresponding to 3.61021 protons on target (POT) was analyzed with improvement to the neutrino flux and interaction modeling compared to previous studies [2]. New photon- and proton-tagged near detector (ND) samples, as well as a new neutrino-enhanced muon-neutrino sample with tagged pions (CC) in the far detector, are included in the analysis for the first time. This letter reports T2K’s neutrino mixing parameter measurements after these improvements.
@@
The T2K experiment— A beam of neutrinos is produced by colliding 30 GeV protons, accelerated by the J-PARC main ring, with a graphite target. These protons interact in the target to produce pions and kaons that are focused into a secondary beam by a system of three magnetic horns and then decay in a 96 m long tunnel. Depending on the direction of the current flowing in the horns, either positively or negatively charged hadrons are focused, producing a neutrino- (-mode) or antineutrino-enhanced (-mode) beam.
@@
Two near detectors are located 280 m away from the graphite target. One is placed on-axis (INGRID) [3] and the other at 2.5∘ off-axis (ND280) with respect to the beam direction. They sample the unoscillated beam by monitoring its direction, intensity and flavor content as well as constraining uncertainties in the neutrino interaction model. The far detector, Super-Kamiokande (SK) [4] is a 50 kton water-Cherenkov detector located beneath a 1 km rock overburden within the Kamioka mine in Japan. It measures the oscillated neutrino flux 295 km away from its production point at off-axis.
@@
The analysis reported in this letter uses all data collected by T2K from January 2010 to February 2020, corresponding to an SK exposure of POT in -mode and POT in -mode. The same POT was used in a previous analysis [2].
@@
Neutrino flux prediction— The neutrino flux Monte Carlo (MC) simulation uses the FLUKA 2011.2 [5, 6] interaction model for proton interactions inside the target and GEANT3 [7] for interactions outside the target. Hadron production in the target is tuned to external datasets [8, 9, 10]. The flux depends on the beamline conditions and the measured proton beam profile, and the prediction is modified to account for conditions during each operating period.
@@
Neutrino interaction modeling— The nominal model prediction was generated with the neutrino event generator NEUT 5.4.0 [12]. Important improvements were made to the estimated neutrino interaction uncertainty for this analysis with the number of free parameters used to describe this uncertainty increasing by 26 from the previous analysis [2] to a total of 75.
@@
Ad hoc uncertainties in the low four-momentum transfer response for CCQE interactions were replaced with parameters that vary the impact of Pauli blocking and the effect of nuclear transparency on the inclusive cross section. An effective parameterization of the latter was derived by comparison to the NuWro [13, 14] implementation of the optical potential [15]. New uncertainties on the nuclear ground state for CCQE events, which parameterize the shell structure of the missing momentum () and missing energy () response in the initial state Spectral Function (SF), were included [16, 17]. The Benhar SF [18] incorporates multi-nucleon knockout in the MeV, MeV/ region. These events account for 5% of all CCQE interactions, a 200% normalization uncertainty was applied to cover differences between NEUT and NuWro predictions. Following Ref. [19], a three-momentum-transfer-dependent freedom on the Nucleon Removal Energy (NRE) was added [20]. Additional freedom for 2p2h interactions was included by separating the existing uncertainties by the target nucleon pair type.
@@
Uncertainties on the free parameters in the NEUT implementation [2] of the Rein–Sehgal SPP model [21] were motivated by fitting [16] to hydrogen- and deuterium-target data [22, 23] from ANL [24, 25] and BNL [26, 27, 28], and checking coverage against hydrocarbon-target data [29, 30, 31, 32, 33, 34, 35]. Variations of these parameters exhibit a minimal change to both the shape of the final-state pion kinematics, which have previously been observed to be poorly predicted at both ND280 and SK [2], and to the relative rate of charged and neutral pion production. To cover observations of pion kinematic spectra in T2K [36] and MINERA [37], new parameters were developed to vary the resonance decay kinematics and the total neutral pion production rate [16]. Previously, NRE effects were only included in NEUT for pionless hard-scatter processes. Here, an approximation of this effect, and an associated uncertainty, was motivated by comparing NEUT predictions to those from NuWro for MeV.
@@
A new 16% uncertainty on the cross section for pion charge-exchange reactions for MeV/ was included [41]. As the new ND selections are sensitive to proton kinematics, an uncertainty was added that varies the fraction of events that include intranuclear nucleon rescattering while keeping the leptonic observables unmodified. The size of this uncertainty (30%) was motivated by a recent analysis of nuclear transparency data [42].
@@
Selected CC events were separated into different samples according to the FGD in which the interaction occurred, the beam mode, the muon charge and the reconstructed final-state particle multiplicities. Negatively-charged muon candidates selected in -mode were divided into five samples per FGD: 1) at least one photon (CC-photon), 2) no photons, pions or protons (CC00p), 3) no photons or pions and at least one proton (CC0Np), 4) no photons and one positively-charged pion (CC1), 5) and all other CC events (CC-Other). Positively and negatively charged muon candidates in -mode are each divided into three samples per FGD: 1) no pions (CC0), 2) one pion with opposite charge to the muon (CC1), 3) and all other CC events (CC-Other). Positively charged muon candidates in -mode were not separately selected as the predicted contamination was 4%, compared to 30% negatively charged muon contamination in -mode [45].
@@
Each sample was binned according to the muon candidate’s momentum, , and the cosine of the angle between the muon momentum vector and the beam direction, . A MC prediction was made using the flux, cross-section and ND280 detector models, and an extended binned likelihood [46] fit to data was performed. Fig. 1 shows the data, prefit and postfit MC distributions for the FGD1 CC00p sample as a function of . The total -value is 0.575, indicating agreement between the data fit result and the input models. The uncertainty on the far detector predicted event rates from cross-section and flux systematics was reduced from approximately 10–15% to 3–4% for each sample by using the ND280 constraint.
@@
Figure 1: Data and model predictions before and after the ND280 fit for the -mode FGD1 CC00p sample binned in . The postfit prediction is broken down by interaction channel. The predicted event rate from contamination in -mode is neglected in the figure as it only contributes at the sub-percent level.
@@
SK event selection—Previous T2K analyses used five SK samples: 1 ring - and -like in both - and -modes and 1 ring -like with 1 decay electron in -mode [2] (tagging a pion below Cherenkov threshold). This analysis introduces a CC1 sample in -mode, which tags pions through two topologies: (1) one ring each from a muon and a charged pion with one or two decay electrons (from the pion and muon decay) and (2) one -like ring (where the charged pion is below Cherenkov threshold) with two decay electrons (from the decay of the muon and the pion). Standard pre-selection cuts common to all SK samples [2] are applied in addition to the particle-identification (PID) cuts for the candidate -like, -like, and decay electron rings. These PID cuts reduce the number of selected background events with primary electrons or neutral pions, and those produced by neutral-current pion production processes.
@@
The addition of the CC1 sample increases the total number of selected -like events by 42.5%, and is affected by different systematics to the dominant 1 ring -like samples. The predicted number of events in this sample at a reference set of oscillation parameter values [47] (, , , eV2/, eV2/, and normal ordering), is 53.5 (116.6) signal (total) events. A total of 134 data events pass the selection criteria with a total systematic uncertainty of 4.3%. The reconstructed neutrino energy for this sample, using only the reconstructed muon information, is
@@
where , , are the proton, muon, and delta baryon rest masses; and , , are the muon-candidate reconstructed energy, three-momentum, and angle with respect to the neutrino beam. Predicted and observed distributions for this sample are presented in Fig. 2.
@@
Figure 2: The distribution for the CC1 -mode SK sample shown for data and MC. The reference oscillation parameter values are described in the main text. The ND280 constraint has been applied to the MC prediction.
@@
Oscillation Analysis— In the analysis, the PDG parameterization of the PMNS matrix [47] was used. Matter effects were included with an average Earth crust density of 2.6 g/cm3. The peak neutrino energy and baseline used by T2K provide sensitivity to the PMNS parameters , , , and the magnitude and sign of the mass-splitting term .
@@
Measurements were carried out with and without a Reactor Constraint (RC). The RC was included as a Gaussian prior on , the PDG world-average [48, 49, 50, 47]. Similarly, for and priors of and were used [47]. Flat priors were used for the other oscillation parameters. The impact of the choice of prior on was investigated and does not alter the conclusions of the analysis.
@@
Two different statistical approaches were used to extract constraints on the oscillation parameters of interest from a likelihood function, the form of which remained unchanged from Ref. [2]. One used a Markov Chain Monte Carlo (MCMC) method to simultaneously fit the data from ND280 and SK, producing posterior distributions from which credible intervals (CI) were constructed. The other performed a piece-wise fit and marginalized over the parameters associated with the propagated ND280 constraint to provide Frequentist confidence intervals, both with the constant and Feldman-Cousins (FC) [51] methods.
@@
Tab. 1 presents the posterior predictive p-values [52, 53, 54] for all FD samples. After accounting for the ‘Look Elsewhere Effect’ using the Bonferroni correction [55], all samples pass the 5% threshold.
@@
Simulated Data Studies (SDS)— These study the impact of alternative interaction models and data-driven modifications at ND280 and SK, to check the completeness of our systematic uncertainty model, and are essential tests of the robustness of our result. Changes to and are deemed significant if the bias to the center of a 2 confidence interval is greater than 50% of the 1 Asimov interval width, or if the size of a one-dimensional, 2, FC-corrected interval changes by more than 10%. If a significant bias is found, additional smearing is applied to the relevant intervals. For , biases to interval boundaries which change the inclusion or exclusion of a physically interesting point in parameter space, e.g. , are reported as part of the result, but due to the non-Gaussian nature of the likelihood additional smearing is not applied to reported intervals.
@@
Various SDS were described in Ref. [2], and have also been carried out in this analysis. New SDS for this analysis are: replacing the default single-pion production model with the Martini et al. 1 model [56, 57] to test the robustness of the model for the CC1 SK sample; altering the default nuclear response to the Hartree-Fock Continuum Random Phase Approximation (HF CRPA) [58, 59] to test the new proton-tagged ND280 samples. In total 19 SDS were performed. New SDS and those with the most significant impacts are shown in Tab. 2. The observed bias on the center of the 2 interval exceeded the bias condition for both the HF CRPA and non-CCQE SDS. As a result, the contour was smeared by , determined from the quadrature sum of the biases on from all SDS. Additionally, a single SDS (SPP low- (Ref. [37])) was found to change the confidence interval of such that is not excluded.
@@
Oscillation results and discussion—Fig. 3 shows credible regions in the – plane produced with and without the RC applied. The measurement of without the RC applied is consistent with the PDG value. The best-fit values, when marginalizing across both mass orderings, with and without the RC applied are and respectively.
The FC-corrected Frequentist confidence intervals for are shown in Fig. 4. The best-fit value was ( ) for normal (inverted) ordering with the RC applied. The data preferred values of close to radians, excluding values around radians at 3, and excluding the CP-conserving values of 0 and at 90% confidence. However, a SDS was found which causes the result to no longer exclude at 90% confidence. The result was statistically limited and can be expected to improve as more data is accumulated.
Figure 4: The change in the best fit observed in the Frequentist analysis when varying and the mass ordering. Shaded regions correspond to the FC-corrected confidence intervals. CP-conserving values are included at the 2 level in normal ordering and excluded at the 3 level in inverted ordering.Figure 5: Frequentist confidence intervals are shown in the – plane, produced using the constant method with the RC applied. The contour smearing to cover SDS as described in the main text is included.
Fig. 5 shows Frequentist confidence intervals in the – plane for both mass orderings. The data exhibited a weak preference for the upper octant of and the normal ordering, with best-fit values of and . The MCMC analysis obtains a Bayes factor with (without) the RC applied of 2.3 (1.4) for the upper octant of over the lower and 2.7 (1.7) for the normal over inverted ordering.
The measurements presented thus far assume the PDG parameterization of the PMNS matrix. The Jarlskog invariant, , is a parameterization-independent way of measuring the scale of CP-violation generated by PMNS oscillations [60]. A zero (non-zero) value for indicates CP-conservation (CP-violation) in three-flavor neutrino mixing. The constraint on , obtained with the MCMC analysis, and the impact of the choice of prior on that constraint, are shown in Fig. 6. The CP-conserving value, , is excluded at the 90% credible for both priors: flat in and flat in . Although changes in the prior were checked, the robustness of these credible intervals in has not been checked with SDS in this analysis.
@@
Conclusions— The T2K collaboration has measured the three-flavor PMNS neutrino oscillation parameters , , , , , and the mass ordering, using POT at SK. The analysis included a new CC1 SK sample, which increased the -mode -like statistics by 42.5%; a new ND event samples in -mode; and significant improvements to the flux and neutrino interaction modeling. The results are consistent with previous analyses [2], showing a weak preference for normal ordering and upper octant of with best-fit values of and . One of the CP conserving values, , is excluded at 90% confidence level and a best-fit value of for normal ordering with the RC applied was found.

### v3

@@
††thanks: also at Université Paris-Saclay††thanks: also at Departament de Fisica de la Universitat Autonoma de Barcelona, Barcelona, Spain††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Moscow Institute of Physics and Technology (MIPT), Moscow region, Russia and National Research Nuclear University “MEPhI”, Moscow, Russia††thanks: also at IPSA-DRII, France††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at JINR, Dubna, Russia††thanks: also at the Graduate University of Science and Technology, Vietnam Academy of Science and Technology††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at BMCC/CUNY, Science Department, New York, New York, U.S.A.††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan
@@
University of Tokyo, Department of Physics, Tokyo, Japan
@@ Abstract
We have made improved measurements of three-flavor neutrino mixing with 19.7(16.3) protons on target in (anti-)neutrino-enhanced beam modes. A new sample of muon-neutrino events with tagged pions has been added at the far detector, as well as new proton and photon-tagged samples at the near detector. Significant improvements have been made to the flux and neutrino interaction modeling. T2K data continue to prefer the normal mass ordering and upper octant of with a near-maximal value of the charge-parity violating phase with best-fit values in the normal ordering of , and eV2.
@@
Introduction— The T2K experiment [1] measures three-flavor neutrino mixing parameters by observing the disappearance of () and the appearance of () over a distance of 295 km in a narrow-band, predominantly or beam, which peaks at an energy of 0.6 GeV. Data corresponding to 3.61021 protons on target (POT) were analyzed with major improvements to the neutrino flux and interaction modeling compared to previous studies [2]. New photon- and proton-tagged near detector (ND) samples, as well as a new, charged current neutrino-enhanced muon-neutrino sample with tagged pions (CC-like) in the far detector, are included in the analysis for the first time. This letter reports T2K’s latest measurements of neutrino mixing parameters after these improvements, which mark a significant step forward in the robustness of our analysis, particularly due to improvements in modeling and constraining neutrino interaction uncertainties, and supersede previous results produced with the same dataset.
Comparable measurements have also been reported by NOvA [3], and future experiments such as Hyper-Kamiokande [4] and DUNE [5] aim to further improve the precision and discovery potential in neutrino oscillation physics.
The T2K experiment—
@@
Protons of 30 GeV energy, accelerated by the J-PARC main ring, collide and interact with a graphite target, producing pions and kaons that are focused by a system of three magnetic horns and decay inside a 96 m long tunnel. Depending on the direction of the current flowing in the horns, either positively or negatively charged hadrons are focused, producing a neutrino- (-mode) or antineutrino-enhanced (-mode) beam.
@@
Two near detectors are located 280 m away from the graphite target — one on-axis (INGRID) [6], and the other 2.5∘ off-axis (ND280) [1] with respect to the beam direction. They sample the unoscillated beam by monitoring its direction, intensity and flavor content as well as constraining uncertainties in the neutrino interaction model. The far detector, Super-Kamiokande (SK) [7], is a 50 kton water-Cherenkov detector located beneath a 1 km rock overburden within the Kamioka mine in Japan. It measures the oscillated neutrino flux 295 km away from its production point at off-axis.
@@
The analysis reported in this letter uses all data collected by T2K from January 2010 to February 2020, corresponding to an SK exposure of POT in -mode and POT in -mode. The same number of POT was used in a previous analysis [2].
@@
Neutrino flux prediction— The neutrino flux Monte Carlo (MC) simulation uses the FLUKA 2011.2 [8, 9] interaction model for proton interactions inside the target and GEANT3 [10] for interactions outside the target. Hadron production in the target is tuned to external datasets [11, 12, 13]. The flux depends on the beamline conditions and the measured proton beam profile. The prediction accounts for the beam conditions during each operating period.
@@
Neutrino interaction modeling— The nominal model prediction was generated with the neutrino event generator NEUT 5.4.0 [15]. In this analysis, important improvements to the estimated neutrino interaction uncertainties were made, increasing the number of free parameters by 26 compared with the previous analysis [2], for a total of 75.
@@
Ad hoc uncertainties in the low four-momentum transfer () response for CCQE interactions were replaced with physics motivated parameters that vary the impact of Pauli blocking and the effect of nuclear transparency on the inclusive cross section. An effective parameterization of the latter was derived by comparison to the NuWro [16, 17] implementation of the optical potential [18]. New uncertainties on the nuclear ground state for CCQE events, which parameterize the shell structure of the missing momentum () and missing energy () response in the initial state Spectral Function (SF), were included [19, 20]. The Benhar SF [21] incorporates multi-nucleon knockout in the MeV, MeV region. These events account for 5% of all CCQE interactions; a 200% normalization uncertainty was applied to cover differences between NEUT and NuWro predictions. Following Ref. [22], a three-momentum-transfer-dependent freedom on the Nucleon Removal Energy (NRE) was added [23]. Additional freedom for 2p2h interactions was included by separating some of the existing uncertainties by the struck nucleon pair (proton–neutron vs. proton–proton/neutron–neutron).
@@
Uncertainties on the free parameters in the NEUT implementation [2] of the Rein–Sehgal SPP model [24] were motivated by fitting [19] to hydrogen- and deuterium-target data [25, 26] from ANL [27, 28] and BNL [29, 30, 31], and checking coverage against hydrocarbon-target data [32, 33, 34, 35, 36, 37, 38]. Variations of these parameters exhibit a minimal change to both the shape of the final-state pion kinematics, which have previously been observed to be poorly predicted at both ND280 and SK [2], and to the relative rate of charged and neutral pion production. To cover measurements of pion kinematic spectra in T2K [39] and MINERA [40], new parameters were developed to vary the resonance decay kinematics and the total neutral pion production rate [19]. Previously, NRE effects were only included in NEUT for pionless hard-scatter processes. Here, an approximation of this effect, and an associated uncertainty, was motivated by comparing NEUT predictions to those from NuWro for MeV.
@@
A new 16% uncertainty on the cross section for pion charge-exchange reactions for MeV was included [44]. As the new ND selections are sensitive to proton kinematics, an uncertainty was added that varies the fraction of events that include intranuclear nucleon rescattering while keeping the leptonic observables unmodified. The size of this uncertainty (30%) was motivated by a recent analysis of nuclear transparency data [45].
@@
Selected CC events were separated into different samples according to the FGD in which the interaction occurred, the beam mode, the muon charge and the reconstructed final-state particle multiplicities. Negatively-charged muon candidates selected in -mode were divided into five samples per FGD: 1) at least one photon (CC-photon); 2) no photons, pions or protons (CC00p); 3) no photons or pions and at least one proton (CC0Np); 4) no photons and one positively-charged pion (CC1); 5) and all other CC events (CC-Other). Positively and negatively charged muon candidates in -mode are each divided into three samples per FGD: 1) no pions (CC0); 2) one pion with opposite charge to the muon (CC1); 3) and all other CC events (CC-Other). Positively charged muon candidates in -mode were not separately selected as the predicted contamination was 4%, compared to 30% negatively charged muon contamination in -mode [48].
Newly added proton-tagged samples offer enhanced abilities to constrain the -dependence of several uncertainties with data. Since ND280 achieves reliable proton tracking efficiency for momenta 450 MeV, samples with (without) protons generally correspond to higher (lower) average . The new photon-tagged sample helps to constrain production uncertainties, and increases the purity of other samples.
@@
Each sample was binned according to the muon candidate’s momentum, , and the cosine of the angle between the muon momentum vector and the beam direction, . A MC prediction was made using the flux, cross-section and ND280 detector models, and an extended binned likelihood [49] fit to data was performed. Fig. 1 shows the data, prefit and postfit MC distributions for the FGD1 CC00p sample as a function of . The total -value is 0.575, indicating agreement between the data fit result and the input models. The uncertainty on the SK predicted event rates from cross-section and flux systematics was reduced from approximately 10–15% to 3–4% for each sample by using the ND280 constraint.
@@
Figure 1: Data and model predictions before and after the ND280 fit for the -mode FGD1 CC00p sample binned in . The fit includes GeV for all ND280 samples, but the range is truncated here for readability. The postfit prediction is broken down by interaction channel. The predicted event rate from contamination in -mode is neglected in the figure as it only contributes at the sub-percent level.
@@
SK event selection—Previous T2K analyses used five SK samples: 1 ring - and -like in both - and -modes and 1 ring -like with 1 decay electron in -mode [2] (tagging a pion below Cherenkov threshold). This analysis introduces a CC1-like sample in -mode, which tags pions through two topologies: (1) one ring each from a muon and a charged pion with one or two decay electrons (from the pion and muon decay); and (2) one -like ring (where the charged pion is below Cherenkov threshold) with two decay electrons (from the decay of the muon and the pion). Standard pre-selection criteria common to all SK samples [2] were applied in addition to the particle-identification (PID) requirements for the candidate -like, -like, and decay electron rings. These PID requirements reduce the number of selected background events with primary electrons or neutral pions, and those produced by neutral-current pion production processes.
@@
The addition of the CC1-like sample increases the total number of selected -like events by 42.5%, although much of the increase is above the oscillation maximum, and is affected by different systematics to the dominant 1 ring -like samples. At a reference set of oscillation parameter values [50] (, , , eV2, eV2, and normal ordering), 53.5 (116.6) signal (total) events were predicted in this sample. A total of 134 data events passed the selection criteria with a total systematic uncertainty of 4.3%. The reconstructed neutrino energy for this sample, using only the reconstructed muon information, is
@@
where , , are the proton, muon, and baryon rest masses; and , , are the muon-candidate reconstructed energy, three-momentum, and angle with respect to the neutrino beam. Predicted and observed distributions for this sample are presented in Fig. 2. Since events do not populate the region of the oscillation dip, we do not expect a significant increase in sensitivity to oscillation parameters. However, this sample provides improved control of background contributions and allows for valuable cross-checks of the cross-section model.
@@
Figure 2: The distribution for the CC1 -mode SK sample shown for data and MC. The fit includes -like (-like) events with (1.25) GeV, but the range is truncated here for readability. The reference oscillation parameter values are described in the main text. The ND280 constraint has been applied to the MC prediction.
@@
Oscillation Analysis— The PDG parameterization of the PMNS matrix [50] was used in the analysis. Matter effects were included with an average Earth crust density of 2.6 g/cm3. The peak neutrino energy and baseline used by T2K provide sensitivity to the PMNS parameters , , , and the magnitude and sign of the mass-splitting term .
@@
Measurements were carried out with and without a Reactor Constraint (RC). The RC was included as a Gaussian prior on , the PDG world-average [51, 52, 53, 50]. Similarly, for and priors of and were used [50]. Flat priors were used for the other oscillation parameters. The impact of the choice of prior on was investigated and does not alter the conclusions of the analysis [54].
@@
Two different statistical approaches were used to extract constraints on the oscillation parameters of interest from a likelihood function, the form of which remained unchanged from Ref. [2]. One used a Markov Chain Monte Carlo (MCMC) method [55, 56], to simultaneously fit the data from ND280 and SK, producing posterior distributions from which credible intervals were constructed. The other performed a piece-wise fit and marginalized over the parameters associated with the propagated ND280 constraint to provide frequentist confidence intervals, both with the constant and Feldman-Cousins (FC) [57] methods.
@@
Tab. 1 presents the posterior predictive p-values [58, 59, 60] for all FD samples. After accounting for the ‘Look Elsewhere Effect’ using the Bonferroni correction [61], all samples pass the 5% threshold. This suggests that the model provides a plausible description of the data within the considered parameter space.
@@
Simulated Data Studies (SDS)— These study the impact of alternative interaction models and data-driven modifications of predictions at ND280 and SK, to check the completeness of our systematic uncertainty model, and are essential tests of the robustness of our result. Changes to and were deemed significant if the bias to the center of a 2 confidence interval was greater than 50% of the 1 Asimov interval width, or if the size of a one-dimensional, 2, FC-corrected interval changed by more than 10%. If a significant bias was found, additional smearing was applied to the relevant intervals. For , biases to interval boundaries which change the inclusion or exclusion of a physically interesting point in parameter space, e.g. , are reported as part of the result; but due to the non-Gaussian nature of the likelihood additional smearing was not applied to reported intervals.
@@
Various SDS were described in Ref. [2], and have also been carried out in this analysis. New SDS for this analysis are: replacing the default single-pion production model with the Martini et al. 1 model [62, 63] to test the robustness of the model for the CC1-like SK sample; altering the default nuclear response to the Hartree-Fock Continuum Random Phase Approximation (HF CRPA) [64, 65] to test the new proton-tagged ND280 samples. In total 19 SDS were performed. New SDS and those with the most significant impact are shown in Tab. 2. The observed bias on the center of the 2 interval exceeded the bias condition for both the HF CRPA and non-CCQE SDS. As a result, the contour was smeared by , determined from the quadrature sum of the biases on from all SDS. Additionally, a single SDS (SPP low- [40]) was found to change the confidence interval of such that is not excluded.
@@
Oscillation results and discussion—Fig. 3 shows credible regions in the – plane produced with and without the RC applied. The measurement of without the RC applied is consistent with the PDG value. When marginalized over both mass orderings, the best-fit value of with (without) the RC applied is ().
The FC-corrected frequentist confidence intervals for are shown in Fig. 4. The best-fit value is ( ) for normal (inverted) ordering with the RC applied. The data prefer values of close to radians, excluding values around radians at 3 in both orderings. The CP-conserving values of 0 and are excluded at the 3 level in inverted ordering. In normal ordering, is excluded at 90% confidence, and although the nominal fit excludes at 90% (Fig. 4), an SDS was found which could move the interval boundary past , so this value is not excluded in the reported result. The result was statistically limited and can be expected to improve as more data is accumulated.
Figure 4: The change in the best fit observed in the frequentist analysis as a function of and the mass ordering. Shaded regions correspond to the FC-corrected confidence intervals.Figure 5: Frequentist confidence intervals are shown in the – plane, produced using the constant method with the RC applied. The contour is smeared to cover the SDS as described in the main text.
Fig. 5 shows frequentist confidence intervals in the – plane for both mass orderings. The data exhibit a weak preference for the upper octant of and the normal ordering, with best-fit values of and . The MCMC analysis obtains a Bayes factor with (without) the RC applied of 2.3 (1.4) for the upper octant of over the lower and 2.7 (1.7) for the normal over inverted ordering.
The measurements presented thus far assume the PDG parameterization of the PMNS matrix. The Jarlskog invariant, , is a parameterization-independent way of measuring the scale of CP-violation generated by PMNS oscillations [66, 67]. A zero (non-zero) value for indicates CP-conservation (CP-violation) in three-flavor neutrino mixing. The constraint on , obtained with the MCMC analysis, and the impact of the choice of prior on that constraint, are shown in Fig. 6. The CP-conserving value, , is excluded at the 90% credible interval for both priors: flat in and flat in . Although changes in the prior were checked, the robustness of these credible intervals in has not been checked with SDS in this analysis.
@@
Conclusions— The T2K collaboration has measured the three-flavor PMNS neutrino oscillation parameters , , , , , and the mass ordering, using POT at SK. The analysis includes a new CC1-like SK sample; new ND event samples in -mode; significant improvements to the flux and neutrino interaction modeling; and a significantly expanded set of SDS to test the robustness of the analysis to out of model changes.
These improvements represent a significant step forward in the development of analysis methods which are robust to mis-modeling issues, a major obstacle for all current and future long-baseline neutrino oscillation experiments.
Our results show a weak preference for normal ordering and the upper octant of with best-fit values of and . One of the CP conserving values, , is excluded at 90% confidence, with a best-fit value of in normal ordering with the RC applied.
A data release including MCMC chains from the fit without the RC applied have been released in Ref. [68].
Acknowledgements—